how many terms of the arithmetic series 10 + 8 + 6 +... will make the sum 24?


1
Expert's answer
2021-07-01T17:26:14-0400

This is the arithmetic sequence with "a_1=10, d=-2."

"S_n=\frac{n}{2}(2a_1+(n-1)d)"

"\frac{n}{2}(20-2(n-1))=24"

"n^2-11n+24=0"

"n=3"